OpenStudy (gabylovesyou):

Let A = {circle, square} and let B = {3, 4, 5}. Find A X B. {circle, square, 3, 4, 5} { } {3, 4, 5} {(circle, 3), (circle, 4), (circle, 5), (square, 3), (square, 4), (square, 5)} Question 21 (Multiple Choice Worth 1 points) [3.06] If f(x) = x^2 - 1, find f(3). 5 2 8 -2

OpenStudy (gabylovesyou):

@AccessDenied @baldymcgee6

OpenStudy (accessdenied):

1) A x B means the set of all ordered pairs whose first entry is an eement of A and second is an element of B. So, you'll have a set looking something like \(\{(a_1, b_2), \ (a_2, b_2)\}\) etc.

OpenStudy (baldymcgee6):

f(x) = x^2 - 1, find f(3) f(3) = (3)^2 -1 f(3) = 9-1 f(3) = 8

OpenStudy (gabylovesyou):

for the first one is it A?

OpenStudy (accessdenied):

No, A) would simply be the union of the sets. This is represented as "A u B." A x B will have ordered pairs as elements of the set, rather than single items from each set.

OpenStudy (gabylovesyou):

it has to be B or C right?

OpenStudy (accessdenied):

It would be D): {(circle, 3), (circle, 4), (circle, 5), (square, 3), (square, 4), (square, 5)} You can see the ordered pairs: (circle, 3), (circle, 4), etc. consist of one element from A first, and then one element from B second. Basically, A x B is like the set of all combinations of elements from A and B. Would that make sense?